OpenStudy (anonymous):

Solve 2x+1/6 = 11x+3/30

OpenStudy (anonymous):

first you need to cross multiple in order to get rid of the denominators

OpenStudy (anonymous):

so then you get 6(11x+3)=30(2x+1)

OpenStudy (anonymous):

once you get that then you would want to multiply them out 66x+18=60x+1

OpenStudy (anonymous):

now you would want to subtract the 60x from 66x 66x+18=60x+1 -60x -60x ________________ 6x+18=1

OpenStudy (anonymous):

now you must subtract the 18 from both sides like i did with the 60x and then you would get 6x=-17

OpenStudy (anonymous):

now divide 6 from both sides and you end up getting x=-17/6 hope this helps let me know if you have any problems

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

your welcome glad to help :)

OpenStudy (anonymous):

I'm a little confused...this is a problem from a review sheet that my teacher gave out, along witht the answers. Her answer to this problem is x=2

OpenStudy (raden):

can u re-type ur question (use draw maybe) ?